OpenStudy (anonymous):

What is the value of x in the solution to the following system of equation x - y = -3 x + 3y = 5

OpenStudy (anonymous):

@beastieman21 Can you help me with this?

OpenStudy (triciaal):

one method by elimination subtract the eqns this will eliminate the x you will have one eqn with one variable. when you solve for y use this value in any of the original eqn to find x, did you get y = 2 x = -1? by substitution from the 1st eqn x = -3 + y use this in the 2nd eqn x + 3y = 5 (-3+y) + 3y = 5 solve for y then use this value of y in x = -3 + y to get x the 3rd method is graphing for each line find the points when x = 0 what is y and when y = 0 what is x this give 2 lines the solution is where they intersect

OpenStudy (triciaal):

any method used will still give x = -1

OpenStudy (anonymous):

x=-1